OpenStudy (anonymous):

use properties of logarithm find the exact value of the expression, log 2^5* log 5^16

OpenStudy (anonymous):

\[\log _{2}5 \times \log _{5}16\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you help me? pls im stuck...

OpenStudy (dumbcow):

are you sure its not \[\log_{}2^{5} * \log_{}5^{16} = 5\log 2*16\log 5 \]

OpenStudy (anonymous):

its wht the cali guy said

OpenStudy (dumbcow):

use change of base rule then \[\log_{2} 5 = \frac{\ln 5}{\ln 2} \] \[\log_{5} 16 = \frac{\ln 16}{\ln 5} \]

OpenStudy (dumbcow):

ln5 will cancel = ln(16)/ln(2)

OpenStudy (anonymous):

tknks

OpenStudy (dumbcow):

you can go even farther too ln(16)/ln(2) = ln(2^4)/ln(2) = 4ln(2)/ln(2) = 4

OpenStudy (anonymous):

4